Find the area of a rectangle of length 16m and perimeter 52m.
Answer by macston(5194)

A=area; L=length=16m; P=perimeter=52m; W=width
P=2(L+W)
52m=2(L+W) Divide each side by 2.
26m=L+W Plug in for L
26m=16m+W Subtract 16m from each side
10m=W
A=L*W
A=16m*10m=160 sq m ANSWER: The area is 160 square meters.